OpenStudy (anonymous):

Consider a drawer full of lose socks (they are not paired together or sorted in manner). Twelve of the socks are navy blue, 6 are brown and 16 are black. If two socks are randomly selected without replacement find the following A.) P (Navy blue then Brown) B. P (Black then Black)?

OpenStudy (mathmagician):

A: (12/34)*(6/33) B: 16*15/(34*33)

OpenStudy (anonymous):

Total socks = 12+6+16 = 34 A. Navy blue = 12/34. Brown = 6/33. Consecutive drawing (12/34)*(6/33). B. (Assuming this is "starting again"): Black = 16/34. Black = 15/33. Consecutive drawing = (16/34)*(15/33).